OpenStudy (anonymous):

Harris purchased a set of golf clubs from the USA for US$900. he had to pay 18% customs duty and 13%HST to import it. calculate the total cost of the clubs. if a similar set of clubs cost C$1000 plus 13% HST, was it cheaper for him to buy it from the USA?, assume that the exchange rate was US$1=C$1.1115

OpenStudy (kash_thesmartguy):

Here, we need to do 2 equations and solve them.

OpenStudy (kash_thesmartguy):

So, what will be the 2 equations? @uju25

OpenStudy (anonymous):

I have no idea

OpenStudy (kash_thesmartguy):

\[900+(900*0.18)+(900*0.13)\]\[900+162+117=1179\] So, 1179 is your answer for the total cost of the first set of clubs. You getting it, right?

OpenStudy (anonymous):

yes, I am

OpenStudy (kash_thesmartguy):

Next, you need to convert C$ into US$. So, do 1000*1.115=?

OpenStudy (anonymous):

115

OpenStudy (anonymous):

1115

OpenStudy (kash_thesmartguy):

Then, \[1115+(1115*0.13)\]\[1115+144.95=$1259.95\]So, is $1259.95 bigger or $1179 bigger?

OpenStudy (kash_thesmartguy):

@uju25

OpenStudy (anonymous):

yes

OpenStudy (kash_thesmartguy):

Yes, what?

OpenStudy (kash_thesmartguy):

The answer is that it wouldn't have been cheaper to buy buy it from Canada. You get it.

OpenStudy (anonymous):

yes, I get it now.. thanks